At 25deg C, the rate constant for the first-order decomposition of a pesticide solution is 6.4 x 10^-3min^-1. If the starting concentration of the pesticide is 0.0314 M, what concentration will remain after 62.0 min at 25deg C?
Answer #1

for first order reaction

K = (2.303 / t) log [A0] / [A]

K = 6.4 x 10-3 min-1

t = 62.0 min

[A0] = 0.0314 M

[A] = ?

6.4 x 10-3 = (2.303 / 62) log [0.0314] / [A]

0.172 = - 1.50 - log [A]

log [A] = - 1.672

[A] = 0.0213 M